OpenStudy (anonymous):

Solve for x: −3|x − 3| = −6 x = 1, x = −1 x = 1, x = 5 x = 0, x = 5 No solutions

OpenStudy (campbell_st):

divide both sides by -3 then | x - 3| = 2 now you need to solve 2 equations x - 3 = 2 or x - 3 = -2

OpenStudy (anonymous):

Oh okay so its the second one. thanks :)

OpenStudy (campbell_st):

yep... looks like it...